Jilk Inc.s contribution margin ratio is 60% and its fixed monthly expenses are $51,000. Assuming that the fixed monthly expeSerfass Corporations contribution format Income statement for July appears below. Sales $369,000 Variable expenses 147, 600
Answer #1

1)Net operating income = Contribution MArgin *sales - fixed cost

= 144000*60%-51,000 = $35,400 (Answer)

2) Degree of Operating leverage = Contibution Margin /opearting Income

=221,400 / 166,050 = 1.33 (Answer)